# and quantum information science

### v2

@@ 1 Preface
The goal of this review is to facilitate the construction of new bridges between different fields of physics that study dynamical phenomena through the lens of quantum complexity. These fields define and use different notions of complexity, which require distinct methods and tools for analysis. Many of these developments are individually reviewed in [1, 2, 3]. We hope to offer a complementary, integrative perspective at the intersection of gravity, quantum information theory, quantum field theory (QFT), and quantum many-body systems. This is a fast moving field and our review captures the state of the art in early 2025.
@@
Historically, the study of quantum complexity originated in the field of quantum computation, generalizing ideas from classical computation. The basic question was, “How difficult is it to solve instances of a particular problem class?” Here, difficulty was measured in terms of a scarce resource like the minimal amount of time, space, or number of circuit gates necessary to implement an algorithm to solve problems exactly, approximately, or even just with high probability. Defined in this way, the concept of complexity functioned as a mathematical tool guiding the design of optimal problem-solving algorithms or circuits, or in the language of quantum mechanics, unitary transformations. Given the arbitrariness inherent in any computational model, the output of complexity theory was usually given in the form of bounds on the minimal amount of resource required to solve general instances of the problem class, as a function of some measure of the problem size. Many related questions have been asked in quantum information theory. For example, how many local operations are required to build a given entangled state of $N$ spins starting with a product state? The answer will depend quantitatively on the “amount” and “kind” of entanglement in the target state, both quantified by information theoretic measures. Local operations are privileged in such analyses because physical processes act locally, and nonlocal coordination or structures must be constructed from local ones. Many of these notions are reviewed in [4].
@@
Over the last two decades, these sorts of questions have also turned out to be relevant for quantum gravity. This is because, in the “holographic” duality between $d+1$-dimensional quantum gravity in Anti-de Sitter (AdS) spacetimes and $d$-dimensional Conformal Field Theories (CFTs), entanglement patterns in the CFT state are geometrically encoded in the dual gravity theory [5]. Matrix Product and tensor network constructions of quantum states provide cartoon models for visualizing the holographic encoding of quantum information, with coarser structure built up deeper in the network and finer structure at the periphery. Questions about the structure, dynamical construction, and time evolution of entangled quantum states can then also be geometrized. We will discuss aspects of the rich literature on this subject as appropriate.
@@
Within this holographic quantum gravity setting, important impetus for thinking about “complexity” as a physical quantity came from the observation that some aspects of geometry on the gravity side of the AdS/CFT duality could not possibly be related to quantum information or entanglement-related quantities in the CFT [6]. Specifically, it turns out that the volumes of spatial sections of black hole spacetimes keep increasing through to times exponential in the black hole entropy, far after the timescale in the dual theory where all entanglement related quantities have equilibrated. Now, if we think of time evolution in a physical theory as a quantum computation enacted by a local Hamiltonian, then the complexity of that computation, understood for example as the size of the minimal local circuit that could perform it, can increase for exponential time. This suggested the idea that some notion of quantum complexity may be geometrized as the volume of spatial section in the AdS gravities dual to CFTs – see Chapter 6 of the review [4] – and seeded many investigations of alternative notions of complexity in physical systems. The resulting ideas have been applied to an enormous range of problems ranging from thermalization in bounded systems, to quantum information in many-body physics, to quantum chaos and integrability, and to the quantization of gravity. Some of these ideas are also shedding light on foundational problems such as the origin of the entropy of black hole horizons, and whether information is lost during black hole evaporation. Some solutions to these problems suggest that black holes appear to have a horizon and to destroy information simply because the fine structure required to decode the information is inaccessible to simple probes (older approaches are discussed in [7] and new developments are reviewed in [8]). The new tools that are now available to define and measure quantum complexity may enable us to study such ideas quantitatively.
@@
Many ideas about complexity in physical systems that are being studied today have antecedents in the classic works of Landauer, Bennet, Zurek and other pioneers of the application of information and computation theory to physics. Interested readers can learn more in the collection [9] where questions like “What is complexity?”, “What causes its increase?”, and “Is there a limit to its increase?” are discussed along with topics of current interest like the onset of chaos and the formation of complex states from simple initial conditions.
@@ 2 Key notions
    where $N$ is the number of degrees of freedom and $t$ is time. $\lambda_{\rm OTOC}$ is defined via semiclassical expansion in a small parameter like $\hbar$ or $1/N$. At the scrambling time $t_{*}\sim\lambda_{\rm OTOC}^{-1}\log(N)$, the first two terms in (1) are equal. Black holes and the large-$N$ SYK model are believed to be examples of fast scramblers, with the largest possible $\lambda_{\rm OTOC}=2\pi/\beta$ [10].
@@
    Jackiw-Teitelboim (JT) gravity: A two-dimensional gravitational theory coupled to a dilaton (a real scalar field $\Phi$[11]. Let $G_{\rm N}$ be Newton’s constant; $\mathcal{M}$, a $2$-dimensional smooth manifold; $g$, the metric’s determinant; $R$, the Ricci scalar; $L$, the AdS radius; $K$, the extrinsic curvature of the induced metric on the manifold’s boundary, $\partial\mathcal{M}$; and $h$, the determinant of that induced metric. The JT-gravity action, excluding topological contributions, is:
[…]
@@
    Out-of-time-order correlator (OTOC): A correlation function measuring the spread of many-body entanglement. Let $S$ be a quantum many-body system, and let $W$ and $V$ denote local unitary or Hermitian operators that act on distant subsystems of $S$. Let $W(t)$ be the Heisenberg-picture time-evolved operator $W$ and let $\rho$ denote a state of $S$. The OTOC [10]111 The OTOC often diverges in QFT. One therefore regularizes it by removing the $\rho$ from the trace’s argument. Instead, most commonly, one places a $\rho^{1/4}$ after each remaining factor. Alternative regularization prescriptions differ in the power of $\rho$ and in operator ordering.
@@
    Quantum chaos: The definition of quantum chaos is a subject of debate. According to an early definition, a quantum-chaotic system has a semiclassical limit that is classically chaotic. Since then, quantum chaos has been defined in terms of nonintegrability, energy-level statistics, out-of-time-ordered correlators, and spectral form factors [12, 13, 14].
@@
    Quantum circuit: A sequence of quantum gates, often represented with time, discretized into layers, running from left to right. Multiple gates can act in each layer (if they commute and if at most one gate acts on each qubit). The number of layers is the circuit’s depth.
@@
    Quantum complexity: The least number of quantum logic gates, selected from a chosen set required to (i) implement a target unitary transformation or (ii) prepare a target state from a given initial one. High-energy theorists sometimes call this quantity quantum computational complexity, which can have a different meaning in computer science. In sections providing  analytic bounds, we will emphasize differences between the exact circuit complexity $\mathcal{C}_{0}(U)$ of a unitary $U$, and the approximate circuit complexity $\mathcal{C}_{\delta}(U)$, the minimal number of gates to approximate $U$ up to error $\delta$. When this distinction is unimportant, we drop the subscript.
@@
    Quantum gate: An elementary unitary operation that acts on a quantum system, usually a set of qubits. A $k$-local quantum gate acts on just $k$ qubits. Example 1-qubit gates include the Hadamard gate; and example 2-qubit gates include the controlled-NOT, or CNOT, gate [15].
@@
    Tensor networks: A graphical representation of quantum states [16] built from multilinear maps called tensors. A tensor’s inputs and outputs, labeled by indices, represent degrees of freedom. Tensor networks are evaluated by contracting tensors by summing over indices. Tensor networks assist variational studies of ground states, and can be used to bound entanglement entropies in a simple, graphical manner. One example, the multiscale entanglement renormalization ansatz (MERA) [17], efficiently approximates ground states of scale-invariant systems. A continuous version (cMERA) approximates ground states of free CFTs [18].
@@
    $|\mathrm{TFD}\rangle\coloneqq\frac{1}{\sqrt{Z_{\beta}}}\sum_{n}e^{-\frac{\beta E_{n}}{2}}|n\rangle|n\rangle\,.$ (5)
    is defined on two copies of the system, where $Z_{\beta}$ is the partition function. The reduced theory on each copy is in a canonical, finite-temperature state.
@@ 3 Dialogue
Sofia: At least two approaches can improve current bounds on complexity. One approach centers on random unitary circuits, which are a toy model for chaotic dynamics. Invoking random unitary circuits, we can apply differential topology and algebraic geometry to manifolds formed by unitary transformations. Also, when considering random unitary circuits, we can use unitary designs—probability distributions that cannot be distinguished from the Haar measure in polynomial time.
The second approach centers on Nielsen’s notion of complexity. Suppose we want to perform some unitary transformation on $n$ qubits. Many circuits implement that transformation. Now observe that the unitary group $\mathrm{SU}(2^{n})$, equipped with a continuous norm, forms a differential manifold. Nielsen considers geodesic paths on this manifold from the identity to the transformation of interest. The geodesic is computed with respect to a “complexity metric” that specifies that distances are small in “simple” directions on the manifold and large in “complex” directions. Here “simple” can be defined in whatever way is appropriate to the problem at hand – for example, we might say that an operation is “simple” if it acts on less than $k$ qubits. The geodesic length in the complexity metric bounds the quantum complexity of the unitary transformation [15, 19]. This bound can be useful, but may not be tight.
@@
Sofia: Certainly! Kolmogorov complexity is the length of the shortest computer program that produces a target output [20]. Another notion, Krylov complexity, quantifies operator growth—an operator’s spread across the space of operators during evolution in the Heisenberg picture [21]. Similarly, spread complexity quantifies the expansion of states across the Hilbert space during time evolution [22].
@@
Sofia: Both have interesting physical applications and are relatively simple to calculate. Krylov complexity was introduced to distinguish chaotic from integrable dynamics of finite-size quantum many-body systems that lack semiclassical limits [21]. Spread complexity relates to spectral form factors, which characterize the energy spectra of many-body quantum-systems [22], and provides tools for quantifying late time dynamics. To calculate the Krylov or spread complexity, we must tridiagonalize the Hamiltonian and there are efficient numerical [21, 22] and analytical [23, 24] techniques for doing so. Therefore, one can calculate these complexities relatively easily.
@@
Sofia: We hope so! Quantum complexity has been proposed to have an Einstein-Rosen bridge’s (ERB’s) volume as a holographic dual. An ERB is a spacetime region that connects a maximally extended Schwarzschild black hole’s two boundaries. In the holographic context, the black hole is in an asymptotically Anti-de Sitter (AdS) spacetime and is dual to the thermofield-double state of a Conformal Field Theory (CFT) on the spacetime boundary. The ERB’s volume was proposed to be a holographic dual to the boundary state’s quantum complexity [6, 25]. More-complicated geometric quantities were conjectured to be dual to quantum complexity, too [26, 27, 28, 29, 30].
@@
More evidence emerges from a class of two-dimensional bulk models. There, a notion of wormhole length was formulated nonperturbatively. This length saturates at times exponentially long in the black hole’s entropy [31] (unlike the classical theory where the length grows forever). This late time behavior mirrors the expected saturation of complexity at times exponentially long in the system size [4]. Moreover, in [32, 33, 34, 35], the authors show that a state’s spread complexity precisely matches the dual wormhole’s classical volume at early times while also saturating at late times, at least in a two-dimensional gravitational setting. What is more, in general dimensions and for any theory of gravity, the saturation of the wormhole’s size at exponential values must be enforced by the finiteness of the black hole’s entropy, as explained in [36, 37].
@@
Sofia: The map I just mentioned—between a wormhole’s volume and spread complexity—hints that spread complexity is appropriate. Yet Refs. [38, 39, 40, 41, 42, 43, 44, 45, 46] point to other notions of complexity. We are not yet certain.
@@
Sagredo: Indeed, entanglement entropy provides a striking example of a holographic mapping that has passed important consistency checks. The dual to a boundary QFT state’s entanglement entropy is the area of a minimal surface in the geometry dual to that state [5]. The dual objects have been identified clearly in this example for various states.
Complexio: But I heard of a paper with the title “Complexity=Anything” which seemed to imply that different notions of complexity could match each of infinitely many gravitational quantities [29, 30, 47]. Every such quantity exhibits linear growth and the switchback effect. But then, I can pick any conjecture I like, and all of them define a geometrical dual to quantum complexity. Does this multiplicity make any sense? Is the cornucopia of such holographic conjectures useful in any way?
@@
Still, I understand the concern – the catchphrase “Complexity = Anything” makes it sound like there is no concrete statement at all. But as we discussed earlier, there is a precise relation between spread complexity and wormhole length at least in the case of two-dimensional gravity. In terms of the “Complexity = Anything” idea, it would help to show that that the ambiguities in alternative definitions of complexity, at least within a single framework like circuit or spread complexity, precisely match the corresponding ambiguities on the gravitational side. It could also be that some symmetries or other desirable properties could restrict the appropriate gravitational quantities to one geometric object, and in this case we would need a principle to select the corresponding notion of complexity. One approach, tried in [48], was to use bulk symplectic forms to relate boundary quantities to circuit complexity.
@@
In the alternative Krylov or spread complexity approaches, we can start with any operator or state. But [22] proved that, for a given initial state, these approaches yield a unique notion of complexity by minimizing the spread of the operator or state over all possible bases describing the evolving system. In the case of spread complexity, different initial states (e.g., random states, thermal states, states localized on some eigenspace) can be used to probe different dynamical properties of the system.

### v3

@@ 1 Preface
The goal of this review is to facilitate the construction of new bridges between different fields of physics that study dynamical phenomena through the lens of quantum complexity. These fields define and use different notions of complexity, which require distinct methods and tools for analysis. Many of these developments are individually reviewed in Susskind:2018pmk, Chapman:2021jbh, Nandy:2024htc. We hope to offer a complementary, integrative perspective at the intersection of gravity, quantum information theory, quantum field theory (QFT), and quantum many-body systems. This is a fast moving field and our review captures the state of the art in early 2025.
@@
Historically, the study of quantum complexity originated in the field of quantum computation, generalizing ideas from classical computation. The basic question was, “How difficult is it to solve instances of a particular problem class?” Here, difficulty was measured in terms of a scarce resource like the minimal amount of time, space, or number of circuit gates necessary to implement an algorithm to solve problems exactly, approximately, or even just with high probability. Defined in this way, the concept of complexity functioned as a mathematical tool guiding the design of optimal problem-solving algorithms or circuits, or in the language of quantum mechanics, unitary transformations. Given the arbitrariness inherent in any computational model, the output of complexity theory was usually given in the form of bounds on the minimal amount of resource required to solve general instances of the problem class, as a function of some measure of the problem size. Many related questions have been asked in quantum information theory. For example, how many local operations are required to build a given entangled state of $N$ spins starting with a product state? The answer will depend quantitatively on the “amount” and “kind” of entanglement in the target state, both quantified by information theoretic measures. Local operations are privileged in such analyses because physical processes act locally, and nonlocal coordination or structures must be constructed from local ones. Many of these notions are reviewed in Aaronson:2016vto.
@@
Over the last two decades, these sorts of questions have also turned out to be relevant for quantum gravity. This is because, in the “holographic” duality between $d+1$-dimensional quantum gravity in Anti-de Sitter (AdS) spacetimes and $d$-dimensional Conformal Field Theories (CFTs), entanglement patterns in the CFT state are geometrically encoded in the dual gravity theory Ryu:2006bv. [^1] While the AdS/CFT is the most well-developed duality between a gravitational and a non-gravitational theory, other proposals have also been discussed in the literature, including different proposals for de Sitter holography (which we refer to in section 7.4) see e.g., the review Galante:2023uyf, celestial holography, see e.g., the review Raclariu:2021zjz and Matrix model of M theory, see e.g., the reviews Brahma:2022ikl, Lin:2025iir among others. Matrix product and tensor network constructions of quantum states provide cartoon models for visualizing the holographic encoding of quantum information, with coarser structure built up deeper in the network and finer structure at the periphery. Questions about the structure, dynamical construction, and time evolution of entangled quantum states can then also be geometrized. We will discuss aspects of the rich literature on this subject as appropriate.
@@
Within this holographic quantum gravity setting, important impetus for thinking about “complexity” as a physical quantity came from the observation that some aspects of geometry on the gravity side of the AdS/CFT duality could not possibly be related to quantum information or entanglement-related quantities in the CFT Susskind:2014moa. Specifically, it turns out that the volumes of spatial sections of black hole spacetimes keep increasing through to times exponential in the black hole entropy, far after the timescale in the dual theory where all entanglement related quantities have equilibrated. Now, if we think of time evolution in a physical theory as a quantum computation enacted by a local Hamiltonian, then the complexity of that computation, understood for example as the size of the minimal local circuit that could perform it, can increase for exponential time. This suggested the idea that some notion of quantum complexity may be geometrized as the volume of spatial section in the AdS gravities dual to CFTs – see Chapter 6 of the review Aaronson:2016vto – and seeded many investigations of alternative notions of complexity in physical systems. The resulting ideas have been applied to an enormous range of problems ranging from thermalization in bounded systems, to quantum information in many-body physics, to quantum chaos and integrability, and to the quantization of gravity. Some of these ideas are also shedding light on foundational problems such as the origin of the entropy of black hole horizons, and whether information is lost during black hole evaporation. Some solutions to these problems suggest that black holes appear to have a horizon and to destroy information simply because the fine structure required to decode the information is inaccessible to simple probes (older approaches are discussed in Balasubramanian:2011dm and new developments are reviewed in Almheiri:2020cfm). The new tools that are now available to define and measure quantum complexity may enable us to study such ideas quantitatively.
@@
Many ideas about complexity in physical systems that are being studied today have antecedents in the classic works of Landauer, Bennet, Zurek and other pioneers of the application of information and computation theory to physics. Interested readers can learn more in the collection zurek2018complexity where questions like “What is complexity?”, “What causes its increase?”, and “Is there a limit to its increase?” are discussed along with topics of current interest like the onset of chaos and the formation of complex states from simple initial conditions.
@@ 2 Key notions
    where $N$ is the number of degrees of freedom and $t$ is time. $\lambda_{\rm OTOC}$ is defined via semiclassical expansion in a small parameter like $\hbar$ or $1/N$. At the scrambling time $t_{*}\sim\lambda_{\rm OTOC}^{-1}\log(N)$, the first two terms in (1) are equal. Black holes and the large-$N$ SYK model are believed to be examples of fast scramblers, with the largest possible $\lambda_{\rm OTOC}=2\pi/\beta$ Maldacena:2015waa.
@@
    Jackiw-Teitelboim (JT) gravity: A two-dimensional gravitational theory coupled to a dilaton (a real scalar field $\Phi$) Mertens:2022irh. Let $G_{\rm N}$ be Newton’s constant; $\mathcal{M}$, a $2$-dimensional smooth manifold; $g$, the metric’s determinant; $R$, the Ricci scalar; $L$, the AdS radius; $K$, the extrinsic curvature of the induced metric on the manifold’s boundary, $\partial\mathcal{M}$; and $h$, the determinant of that induced metric. The JT-gravity action, excluding topological contributions, is:
    $S_{\rm JT}[g,\Phi]\coloneqq\frac{1}{16\pi G_{\rm N}}\int_{\mathcal{M}}d^{2}x\,\sqrt{-g}\>\Phi\left(R+\frac{2}{L^{2}}\right)+\frac{1}{8\pi G_{\rm N}}\int_{\partial\mathcal{M}}dx\,\sqrt{-h}\>\Phi\left(K-\frac{1}{L}\right)\,.$ (2)
@@
    Out-of-time-order correlator (OTOC): A correlation function measuring the spread of many-body entanglement. Let $S$ be a quantum many-body system, and let $W$ and $V$ denote local unitary or Hermitian operators that act on distant subsystems of $S$. Let $W(t)$ be the Heisenberg-picture time-evolved operator $W$ and let $\rho$ denote a state of $S$. The OTOC Maldacena:2015waa [^2] The OTOC often diverges in QFT. One therefore regularizes it by removing the $\rho$ from the trace’s argument. Instead, most commonly, one places a $\rho^{1/4}$ after each remaining factor. Alternative regularization prescriptions differ in the power of $\rho$ and in operator ordering.
@@
    Quantum chaos: The definition of quantum chaos is a subject of debate. According to an early definition, a quantum-chaotic system has a semiclassical limit that is classically chaotic. Since then, quantum chaos has been defined in terms of nonintegrability, energy-level statistics, out-of-time-ordered correlators, spectral form factors and adiabatic gauge potential Haake_10_Quantum, DAlessio_16_From, Gogolin_16_Equilibration, Pandey:2020tru. In the remainder of this review, we will argue that Nielsen (Sec. 5) and Krylov/spread (Sec. 6) complexities can also be used as diagnostic tools of chaos.
@@
    Quantum circuit: A sequence of quantum gates (see key notion 14), often represented with time, discretized into layers, running from left to right. Multiple gates can act in each layer (if they commute and if at most one gate acts on each qubit). The number of layers is the circuit’s depth.
@@
    Quantum complexity: The least number of quantum gates (see key notion 14 for the definition of quantum gate), selected from a chosen set required to (i) implement a target unitary transformation or (ii) prepare a target state from a given initial one. High-energy theorists sometimes call this quantity quantum computational complexity, which can have a different meaning in computer science. In sections providing analytic bounds, we will emphasize differences between the exact circuit complexity $\mathcal{C}_{0}(U)$ of a unitary $U$, and the approximate circuit complexity $\mathcal{C}_{\delta}(U)$, the minimal number of gates to approximate $U$ up to error $\delta$. When this distinction is unimportant, we drop the subscript.
@@
    Quantum gate: An elementary unitary operation that acts on a quantum system, usually a set of qubits. A $k$-local quantum gate acts on just $k$ qubits. Example 1-qubit gates include the Hadamard gate; and example 2-qubit gates include the controlled-NOT, or CNOT, gate nielsen2010quantum.
@@
    Tensor networks: A graphical representation of quantum states Cirac:2020obd built from multilinear maps called tensors. A tensor’s inputs and outputs, labeled by indices, represent degrees of freedom. Tensor networks are evaluated by contracting tensors by summing over indices. Tensor networks assist variational studies of ground states, and can be used to bound entanglement entropies in a simple, graphical manner. One example, the multiscale entanglement renormalization ansatz (MERA) Vidal:2007hda, efficiently approximates ground states of scale-invariant systems. A continuous version (cMERA) approximates ground states of free CFTs Haegeman:2011uy.
@@
    $|\mathrm{TFD}\rangle\coloneqq\frac{1}{\sqrt{Z_{\beta}}}\sum_{n}e^{-\frac{\beta E_{n}}{2}}|n\rangle_{1}|n^{*}\rangle_{2}\,$, (5)
    is defined on two copies of the system, where $Z_{\beta}$ is the partition function. In the last equation we defined $|n^{*}\rangle\equiv\Theta|n\rangle$ with $\Theta$ an anti-unitary operator, such as CPT, that commutes with the Hamiltonian, see e.g., Cottrell:2018ash. The conjugation is sometimes kept implicit in order to simplify the notation. The reduced theory on each copy is in a canonical, finite-temperature state.
@@ 3 Dialogue
Sofia: At least two approaches can improve current bounds on complexity. One applies to random unitary circuits (see Sec. 4.3-4.5), which are a toy model for chaotic dynamics. Invoking random unitary circuits, we can apply differential topology and algebraic geometry to manifolds formed by unitary transformations. Also, when considering random unitary circuits, we can use unitary designs—probability distributions that cannot be distinguished from the Haar measure in polynomial time.
The second approach (see Sec. 5) centers on Nielsen’s notion of complexity. Suppose we want to perform some unitary transformation on $n$ qubits. Many circuits implement that transformation. Now observe that the unitary group $\mathrm{SU}(2^{n})$, equipped with a continuous norm, forms a differential manifold. Nielsen considers geodesic paths on this manifold from the identity to the transformation of interest. The geodesic is computed with respect to a “complexity metric” that specifies that distances are small in “simple” directions on the manifold and large in “complex” directions. Here “simple” can be defined in whatever way is appropriate to the problem at hand – for example, we might say that an operation is “simple” if it acts on less than $k$ qubits. The geodesic length in the complexity metric bounds the quantum complexity of the unitary transformation nielsen2010quantum, Brown:2022phc. This bound can be useful, but may not be tight.
@@
Sofia: Certainly! Kolmogorov complexity is the length of the shortest computer program that produces a target output Kolmogorov. Another notion, Krylov complexity, quantifies operator growth—an operator’s spread across the space of operators during evolution in the Heisenberg picture Parker:2018yvk. Similarly, spread complexity quantifies the expansion of states across the Hilbert space during time evolution Balasubramanian:2022tpr.
@@
Sofia: Both have interesting physical applications and are relatively simple to calculate. Krylov complexity was introduced to distinguish chaotic from integrable dynamics of finite-size quantum many-body systems that lack semiclassical limits Parker:2018yvk. Spread complexity relates to spectral form factors, which characterize the energy spectra of many-body quantum-systems Balasubramanian:2022tpr, and provides tools for quantifying late time dynamics. To calculate the Krylov or spread complexity, we must tridiagonalize the Hamiltonian and there are efficient numerical Parker:2018yvk, Balasubramanian:2022tpr and analytical Balasubramanian:2022dnj, Balasubramanian:2023kwd techniques for doing so. Therefore, one can calculate these complexities relatively easily.
@@
Sofia: We hope so! Quantum complexity has been proposed to have an Einstein-Rosen bridge’s (ERB’s) volume as a holographic dual. An ERB is a spacetime region that connects a maximally extended Schwarzschild black hole’s two boundaries. In the holographic context, the black hole is in an asymptotically Anti-de Sitter (AdS) spacetime and is dual to the thermofield-double state of a Conformal Field Theory (CFT) on the spacetime boundary. The ERB’s volume was proposed to be a holographic dual to the boundary state’s quantum complexity Susskind:2014moa, Stanford:2014jda. More-complicated geometric quantities were conjectured to be dual to quantum complexity, too Brown:2015bva, Brown:2015lvg, Couch:2016exn, Belin:2021bga, Belin:2022xmt.
@@
More evidence emerges from a class of two-dimensional bulk models. There, a notion of wormhole length was formulated nonperturbatively. This length saturates at times exponentially long in the black hole’s entropy Iliesiu:2021ari (unlike the classical theory where the length grows forever). This late time behavior mirrors the expected saturation of complexity at times exponentially long in the system size Aaronson:2016vto. Moreover, in Berkooz:2018qkz, Lin:2022rbf, Rabinovici:2023yex, Balasubramanian:2024lqk, the authors show that a state’s spread complexity precisely matches the dual wormhole’s classical volume at early times while also saturating at late times, at least in a two-dimensional gravitational setting. What is more, in general dimensions and for any theory of gravity, the saturation of the wormhole’s size at exponential values must be enforced by the finiteness of the black hole’s entropy, as explained in Balasubramanian:2022gmo, Balasubramanian:2022lnw.
@@
Sofia: The map I just mentioned—between a wormhole’s volume and spread complexity—hints that spread complexity is appropriate. Yet Refs. Caputa:2017urj, Caputa:2017yrh, Caputa:2018kdj, Chandra:2021kdv, Chagnet:2021uvi, Chandra:2022pgl, Erdmenger:2022lov, Arean:2024pzo, Caceres:2024edr point to other notions of complexity. We are not yet certain.
@@
Sagredo: Indeed, entanglement entropy provides a striking example of a holographic mapping that has passed important consistency checks. The dual to a boundary QFT state’s entanglement entropy is the area of a minimal surface in the geometry dual to that state Ryu:2006bv. The dual objects have been identified clearly in this example for various states.
Complexio: But I heard of a paper with the title “Complexity=Anything” which seemed to imply that different notions of complexity could match each of infinitely many gravitational quantities Belin:2021bga, Belin:2022xmt, Jorstad:2023kmq. Every such quantity exhibits linear growth and the switchback effect. But then, I can pick any conjecture I like, and all of them define a geometrical dual to quantum complexity. Does this multiplicity make any sense? Is the cornucopia of such holographic conjectures useful in any way?
@@
Still, I understand the concern – the catchphrase “Complexity = Anything” makes it sound like there is no concrete statement at all. But as we discussed earlier, there is a precise relation between spread complexity and wormhole length at least in the case of two-dimensional gravity. In terms of the “Complexity = Anything” idea, it would help to show that that the ambiguities in alternative definitions of complexity, at least within a single framework like circuit or spread complexity, precisely match the corresponding ambiguities on the gravitational side. It could also be that some symmetries or other desirable properties could restrict the appropriate gravitational quantities to one geometric object, and in this case we would need a principle to select the corresponding notion of complexity. One approach, tried in Belin:2018bpg, was to use bulk symplectic forms to relate boundary quantities to circuit complexity.
@@
In the alternative Krylov or spread complexity approaches, we can start with any operator or state. But Balasubramanian:2022tpr proved that, for a given initial state, these approaches yield a unique notion of complexity by minimizing the spread of the operator or state over all possible bases describing the evolving system. In the case of spread complexity, different initial states (e.g., random states, thermal states, states localized on some eigenspace) can be used to probe different dynamical properties of the system.